# and quantum information science

### v1

@@ 7.6 Computational pseudorandomness and the holographic dictionary
In all the above discussions, the protagonist was the wormhole’s volume. It was observed by Susskind [206] that the steady growth of a wormhole’s volume poses a challenge in the context of the AdS/CFT correspondence: The dual quantity to the wormhole’s volume cannot be a local observable as those are expected to equilibrate and hence saturate quickly in a quantum theory. Instead, it was proposed that the dual of the wormhole’s volume is a global property of the dual state such as quantum complexity. However, from a complexity-theoretic point of view, there appears to be a mismatch between these two quantities. The volume of a wormhole is a comparably simple quantity that can be estimated from a wormholes description. In comparison, quantum complexity is notoriously difficult to compute or even bound. Making this mismatch concrete, Ref. [345] argues that either quantum gravity cannot be simulated efficiently on a quantum computer, or the holographic dictionary requires exponentially long computations. The former would violate the extended Church-Turing thesis– the widely held belief that all physical processes can be efficiently simulated on a quantum computer. More precisely, Ref. [345] models the dual CFT as a quantum system initialized in a thermofield-double state $|\mathrm{TFD}\rangle$ dual to a zero-volume wormhole. This state is undergoing time evolution under a fixed Hamiltonian $H_{\mathrm{CFT}}$ such that $|\mathrm{TFD}\rangle$ is a low-energy state for $H_{\mathrm{CFT}}$. However, in this time evolution, the system undergoes a number of random ‘shocks’ acting locally on single qubits:
@@
In Ref. [345], the authors conjecture that the time-evolution with shocks in Eq. (129), while efficiently simulable on a quantum computer, is computationally indistinguishable from a Haar random state. Such an asymmetry between the easiness of preparing a state and the hardness of distinguishing it from a random state is called quantum pseudorandomness [346, 347]. More precisely, an ensemble of states, parameterized by a random key $k$ is called pseudorandom if 1) it can be efficiently prepared by polynomial-sized quantum circuits and 2) any quantum computer requires superpolynomial time to distinguish the ensemble from Haar random states. This implies that even if the CFT evolution under shocks is computationally simple, its output states may appear indistinguishable from truly random states, making certain bulk properties computationally inaccessible from the CFT side. In particular, quantum circuit complexity cannot be computable using only polynomial-time quantum computations and measurements as such a distinguisher would break pseudorandomness. The existence of pseudorandom states was proven in Ref. [346] under mild cryptographic assumptions.
@@
We can now consider two time evolutions as in Eq. (129) for vastly different times, both undergoing shocks. Pseudorandomness implies that the two resulting states cannot be efficiently distinguished from Haar randomness and, therefore, from each other. However, the two states will correspond to wormholes of different volumes in any holographic dictionary, and the volume is efficiently computable from the AdS state using only coarse-grained properties of the metric. This contradiction, it is argued in Ref. [345], can only be resolved if 1) quantum gravity is not efficiently simulable on a quantum computer or 2) the holographic dictionary cannot be computed in polynomial time. Notice that this argument actually does not require the Volume$=$Complexity conjecture. In fact, the existence of pseudorandomness implies that any quantity that distinguishes between deep time-evolutions under shocks and short time evolutions under shocks needs to be computationally hard to estimate.
@@
A central concept in the AdS/CFT correspondence is the duality of entanglement and geometry as explained by the Ryu-Takayanagi formula [337]. It was shown in Ref. [348] that “geometric reconstruction” is hard even for states whose entanglement entropy satisfies inequalities refered to as the holographic entropy cone [349]. These inequalities define a cone in the space spanned by the entanglement entropies of subregions. In particular, satisfying the holographic entropy cone is a necessary requirement for the Ryu-Takayanagi formula. Geometric reconstruction refers to the task of estimating basic properties of the underlying bulk geometry from a CFT state (see, e.g., [350, 351, 352, 353, 354]). Similar to the argument made in Ref. [345] this hardness result uses pseudorandomness and, more precisely, the concept of pseudoentanglement. Pseudoentanglement [355, 356] refers to the phenomenon that even exponentially large gaps in entanglement can be computationally hard to detect [355], which indicates that entanglement is an “unfeelable” quantity, similar to quantum circuit complexity. In particular, Ref. [348] uses recently constructed low-entangling pseudorandom unitaries [357] to show that two states with arbitrary bulk geometries can correspond to indistinguishable states in the CFT. As the construction in Ref. [348] does not require event horizons, the result suggests that reconstruction of the bulk geometry might even be hard outside the event horizon. Previously, the presence of an event horizon was identified by Susskind [358] as a way to reconcile the Church-Turing thesis with the duality of volume and complexity: An observer that needs to cross the event horizon to estimate the volume will then be unable to communicate the findings to an outside observer. The latter, it was argued, would be in agreement with an exponentially hard-to-estimate dual quantity, such as complexity. The construction in [348] does not take the CFT Hamiltonian into account and produces high-energy states. Consequently, more work is required to draw conclusions about realistic holographic scenarios.
@@
These results raise questions about the broader landscape of computational complexity in holography, particularly in relation to the Python’s Lunch conjecture [359]. This conjecture posits that if a geometry contains a local minimal surface that is not globally minimal, then finding the CFT operator corresponding to a local bulk operator inserted in the region surrounded by the local minimal surface is exponentially hard. The conjecture suggests that if the bulk contains locally minimal but not globally minimal surfaces, then efficient application of operators in the CFT is exponentially complex. The strong Python’s lunch conjecture then further posits that locally minimal surfaces are the only obstruction to efficient “operator reconstruction” [348]. A follow-up [360] shows that the strong Python’s lunch conjecture is compatible with the findings of Ref. [345] on the reconstruction of a wormhole’s volume. The results of Ref. [348] then suggest that the complexity of operator reconstruction in the CFT does not always align with the complexity of bulk geometry reconstruction. These questions about the complexity of the holographic dictionary, inspired by the volume=complexity, require and deserve future investigations.
@@ 8 Paradigms for complexity III: tensor-network-inspired definitions of complexity
Path-integral complexity [28, 29] was the first proposal for complexity in QFTs motivated by holography. In the early days of holographic complexity, the concept of complexity was often discussed in the context of tensor networks, which provide variational ansätze for many-body quantum wavefunctions. In particular, optimized tensor networks such as, e.g., MERA [12, 361] or its continuous counterpart cMERA [13, 83], give rise to a geometric structure that reflects the entanglement patterns of a quantum state.
This picture strongly resonated with the earlier idea of holography emerging from tensor networks [4]. This proposal started a new line of research, termed AdS/TN (Anti-de Sitter/Tensor Network). In fact, it was within the AdS/TN framework that Susskind [18] argued that the growth of the Einstein-Rosen bridge should be related to complexity, estimated by the number of tensors (the effective volume of the network) in the tensor-network representation of the time-evolved thermofield-double state of Hartman and Maldacena [204]. In Sec. 8.1, we review a definition of a notion of complexity in CFT, inspired by the above-mentioned parallelism with tensor networks, called path-integral complexity. We discuss the holographic interpretation of this quantity in a bulk AdS spacetime in Sec. 8.2.
@@ 8.1 Complexity from path-integral optimization
Motivated by these developments, Ref. [28] aimed to extend the tensor-network intuitions to general, strongly-interacting QFTs and define a natural measure of complexity using path integrals. A brief review of their construction proceeds as follows.
In QFTs, wave functions are prepared using path integrals over physical space and Euclidean time. They are computed by integrating over all field configurations, subject to specified boundary conditions on the fields. As an example, consider two-dimensional CFTs on a flat, Euclidean plane $\mathbb{R}^{2}$ with coordinates $(\tau,x)$, and collectively denote all fields in the CFT by $\Phi(\tau,x)$. The ground state wave function $\Psi_{\rm CFT}[\Phi(x)]$, which in QFTs becomes a functional of the boundary conditions imposed at the time slice $\tau=0$: $\Phi(x)\coloneqq\Phi(x,\tau=0)$, is defined as
@@
where $S_{\rm CFT}$ is the Euclidean action of the 2D CFT and $D\tilde{\Phi}$ is the path integrals measure. [^29] Formally one may need to deform the theory by putting it in a finite volume or introducing a small mass in order to have a gapped spectrum so that the path integral (130) acts as a proper vacuum projector.
The main new idea in [28] was to introduce a non-trivial background metric on the space where the path integral (130) is performed (keeping the boundary conditions for the fields fixed), and interpret this metric as an (unoptimized) density of gates of a continuous tensor network. For our 2D CFT example, all the metrics can be put into the Weyl-flat form parametrized by a scalar function $\phi(\tau,x)$ as follows:
@@
Figure 13: A cartoon explaining path-integral optimization as an analog of tensor-network optimization in CFTs. Figure from [28] with Euclidean time denoted as $z=-\tau$.
Next, one can introduce the actual optimization procedure. Namely, if we now compare the wave functions computed with a background flat metric (denoted with $\Psi_{\text{CFT}}^{\phi_{0}}[\Phi(x)]$) and with a curved metric (denoted with $\Psi_{\text{CFT}}^{\phi}[\Phi(x)]$), they are proportional to each other
@@
where $c$ is the central charge of the 2D CFT. [^30] The coefficient $\mu$ of the potential term can be set to $1$ by shifting $\phi$. Intuitively, the proportionality factor in Eq. (132) can be interpreted as a measure of how many repetitions of path-integration per site we perform to construct the state. Hence, the optimal path integral is selected by minimizing the Liouville action with respect to $\phi$, i.e., picking the metric that solves the Liouville equation
@@
with the boundary condition (131). Recall that the metric encoded the density of gates in the tensor network. This procedure is therefore called the path-integral optimization.
For example, for the CFT vacuum that is obtained by performing a Euclidean path integral on the half-plane, the optimal geometry is given by the hyperbolic plane. Similar results can be obtained for path integrals on the circle, that prepares the vacuum state of a CFT on the circle; as well as path integrals on the strip of size $\beta$ used to prepare the thermofield-double state with inverse temperature $\beta$. In the former case, the background obtained via the path-integral optimization procedure is the hyperbolic geometry of the Poincaré disc. In the latter case, the background geometry is the hyperbolic strip, also called the Euclidean trumpet geometry. We can think about these geometries as the counterparts of effective tensor-network geometry that emerges after the optimization procedure. In fact, in the spirit of AdS/TN discussions in holographic CFTs, it is tempting to interpret all these solutions as geometries on time slices of higher-dimensional AdS spacetime. We will return to this point in Sec. 8.2.
Another central idea of references [28, 29] was to interpret the Liouville action as a novel measure of complexity for the wave functions prepared using Euclidean path integrals, dubbed path-integral complexity. The main reason to advocate this interpretation was that the proportionality of the on-shell Liouville action to the volume of the space on which the path integral was computed. Indeed, after fixing the normalization such that the integral of the Weyl factor over a square region of the size $\epsilon$ has value 1, we can interpret the volume of the Euclidean space as the number of tensors in the optimal network. The on-shell action associated with the CFT vacuum state reads
@@
This result is also consistent with qualitative expectations from holography, that we discuss in the next subsection.
@@
Later on, researchers extended the path-integral optimization method to other settings. References [28, 29] proposed a higher-dimensional generalization of the action (135). The optimization of 2D CFTs deformed by relevant operators was studied in [362], where a modified Liouville action was minimized over both metrics and local deformation couplings. The authors of [363] investigated path integrals for inhomogeneous CFTs, interpreted as CFTs coupled to non-trivial background metrics. Soon after [28], reference [364] argued that a variation of path integral complexity in holographic 2D CFTs leads to Einstein’s equations in the bulk of AdS3. Moreover, the Liouville action (or its covariant version, called the Polyakov action) motivated the treatment of Nielsen’s complexity by introducing trajectories on the unitary manifold generated by sequences of 2D conformal transformations [30] (see Sec. 5.5.2).
@@
Liouville action also inspired a series of developments in the tensor-network community. They used the path integral approach to cMERA to define tensor networks in terms of the background geometry on which the state is prepared [365, 366]. In particular, they showed how the background metric’s parameters translate into the tensor network’s details. This was later used in [367] to argue for a Nielsen-type interpretation of the Liouville action in terms of an appropriate choice of the cost function on the complexity geometry. As opposed to the developments on Nielsen’s complexity discussed in Sec. 5, the work [367] utilized necessarily circuits that contained both unitary and Hermitian gates associated with components of the energy-momentum tensor operator. This is similar in spirit to the results [320] discussed in Sec. 7.5.1 and has the same origin: Euclidean time evolution is an efficient and common theoretical way of preparing relevant states. This points towards a mathematical generalization of the Nielsen’s complexity that involves both unitary and Hermitian exponents of the generators.
@@
Finally, the results in Ref. [367] pointed out that a general cost function for path integral optimization will also contain positive powers of the cut-off $\epsilon$ associated with higher derivatives of $\phi$ in the spirit of an effective field theory. However, the solutions of Eq. (134) acquire derivatives that are themselves of the order of $1/\epsilon$. This breaks the hierarchy within these more general cost functions that left the contribution (133) as the dominant term and calls for a systematic inclusion of the UV cut-off $\epsilon$. This moves us directly to the subject of the next section.

### v2

@@ 7.7 Computational pseudorandomness and the holographic dictionary
In all the above discussions, we mainly considered the volume of wormholes as a dual description of some notion of the complexity of time-evolving thermofield double state. It was observed by Susskind [247] that the steady growth of a wormhole’s volume poses a challenge in the context of the AdS/CFT correspondence: The dual quantity to the wormhole’s volume cannot be a local observable as those are expected to equilibrate and hence saturate quickly in a quantum theory. Instead, [247] proposed that the dual of the wormhole’s volume is a global property of the dual state such as quantum complexity. However, from a complexity-theoretic point of view, there appears to be a mismatch between these two quantities. The volume of a wormhole is a simple quantity that can be easily calculated from the gravitational description. In comparison, quantum complexity is notoriously difficult to compute or even bound. Making this mismatch concrete, Ref. [397] argues that either quantum gravity cannot be simulated efficiently on a quantum computer, or some entries in the holographic dictionary require exponentially long computations. The former would violate the extended Church-Turing thesis – the widely held belief that all physical processes can be efficiently simulated on a quantum computer. More precisely, Ref. [397] models the dual CFT as a quantum system initialized in a thermofield-double state $|\mathrm{TFD}\rangle$ dual to a zero-volume wormhole. This state is undergoing time evolution under a fixed Hamiltonian $H_{\mathrm{CFT}}$ such that $|\mathrm{TFD}\rangle$ is a low-energy state for $H_{\mathrm{CFT}}$. However, in this time evolution, the system undergoes a number of random ‘shocks’ acting locally on single qubits:
@@
In Ref. [397], the authors conjecture that the time-evolution with shocks in Eq. (142), while possible to efficiently simulate on a quantum computer, is computationally indistinguishable from a Haar random state. Such an asymmetry between the easiness of preparing a state and the hardness of distinguishing it from a random state is called quantum pseudorandomness [398, 399]. More precisely, an ensemble of states, parameterized by a random key $k$ is called pseudorandom if 1) it can be efficiently prepared by polynomial-sized quantum circuits and 2) any quantum computer requires superpolynomial time to distinguish the ensemble from Haar random states. This implies that even if the CFT evolution under shocks is computationally simple, its output states may appear indistinguishable from truly random states, making certain bulk properties computationally inaccessible from the CFT side. In particular, quantum circuit complexity cannot be computable using only polynomial-time quantum computations and measurements, as such computations would be able to rapidly distinguish pseudorandom and random states, which should be impossible. The existence of pseudorandom states was proven in Ref. [398] under mild cryptographic assumptions.
@@
We can now consider two time evolutions as in Eq. (142) for vastly different times, both undergoing shocks. Pseudorandomness implies that the two resulting states cannot be efficiently distinguished from Haar random states and, therefore, from each other. However, the two states will correspond to wormholes of different volumes in any holographic dictionary, and the volume is efficiently computable from the AdS state using only coarse-grained properties of the metric. Ref. [397] argues that this tension can only be resolved if either 1) quantum gravity cannot be efficiently simulated on a quantum computerm or 2) the holographic dictionary cannot be computed in polynomial time. Notice that this argument actually does not require the Volume$=$Complexity conjecture. In fact, the existence of pseudorandomness implies that any quantity that distinguishes between long and short time evolutions under shocks must be computationally hard to estimate.
@@
A central concept in the AdS/CFT correspondence is the relation between entanglement entropy of subregions in the CFT and the area of a corresponding extremal surface in AdS space, as specified by the Ryu-Takayanagi formula [5]. The existence of this formula suggests that there may be a way of reconstructing AdS geometry from patterns of entanglement in the CFT – a task that we may call geometric reconstruction (see, e.g., [400, 401, 402, 403] for earlier approach to reconstructing local operators). With this in mind, suppose that we consider CFT states that satisfy the restrictive holographic entropy cone inequalities [404]. These inequalities define a cone in the space spanned by the entanglement entropies of subregions and are a necessary requirement for geometrization of entanglement by the Ryu-Takayanagi formula. It was shown in Ref. [405] that “geometric reconstruction” is hard even for states whose entanglement entropy satisfies the holographic entropy cone. Similar to the argument made in Ref. [397] this hardness result uses pseudorandomness and, more precisely, the concept of pseudoentanglement. Pseudoentanglement [406, 407] refers to the phenomenon that even exponentially large gaps in entanglement can be computationally hard to detect [406], which indicates that entanglement is an “unfeelable” quantity, similar to quantum circuit complexity. In particular, Ref. [405] uses recently constructed low-entangling pseudorandom unitaries [408] to show that two states with arbitrarily different bulk geometries can correspond to indistinguishable states in the CFT. As the construction in Ref. [405] does not require event horizons, the result suggests that reconstruction of the bulk geometry might even be hard outside the event horizon. Previously, the presence of an event horizon was identified [409] as a way to reconcile the Church-Turing thesis with the duality of volume and complexity: an observer that needs to cross the event horizon to estimate the volume will then be unable to communicate the findings to an outside observer. This would be consistent with an exponentially hard-to-estimate dual quantity, such as complexity. Likewise, the authors of [236] argued that when the time evolution operator of a holographic CFT is approximately pseudorandom (or Haar random) on some code subspace, then there must be an event horizon in the corresponding bulk dual. The apparent tension with [405] may be resolved by the fact that the construction in the latter does not take the CFT Hamiltonian into account and produces high-energy states.
@@
These results raise questions about the broader landscape of computational complexity in holography, particularly in relation to the Python’s Lunch conjecture [410]. This conjecture posits that if a geometry contains a local minimal surface that is not globally minimal, then finding the CFT operator corresponding to a local bulk operator inserted in the region surrounded by the local minimal surface is exponentially hard. The conjecture suggests that if the bulk contains locally minimal but not globally minimal surfaces, then efficient application of operators in the CFT is exponentially complex. The strong Python’s lunch conjecture then further posits that locally minimal surfaces are the only obstruction to efficient “operator reconstruction” [405]. A follow-up [231] shows that the strong Python’s lunch conjecture is compatible with the findings of Ref. [397] on the reconstruction of a wormhole’s volume. Similarly, [411, 412] show that it is exponentially hard to decode the state of a field in the interior of a black hole from the distant Hawking radiation entangled with it, as expected from the Hayden-Preskill scenario [413]. The results of Ref. [405] then suggest that the complexity of operator reconstruction in the CFT does not always align with the complexity of bulk geometry reconstruction. These questions about the complexity of the holographic dictionary, inspired by the volume=complexity, require and deserve future investigations.
Finally, recent developments relate the growth of the volume of Einstein-Rosen bridges to the randomness of the underlying microscopic quantum gravity state [90, 91]. Concretely, the authors of [90, 91] found gravitational duals to continuous random circuits (see Sec. 4.5 for a survey of random circuits in the context of circuit complexity). The resulting geometries are long wormholes with matter inhomogeneities, dubbed ER-caterpillars. The construction allowed the authors to demonstrate a relation between the length of the ER bridge and the degree of microscopic randomness of the state as a function of time. More precisely, [90, 91] showed that an ER-caterpillar of length $l$ has an underlying quantum state well-approximated by a quantum state k-design with $k\propto l$ when $l$ is large. It would be interesting to understand the implications of this construction in the context of the pseudorandomness mentioned above.
@@ 8 Paradigms for complexity III: tensor-network-inspired definitions of complexity
Path-integral complexity [38, 39] was the first approach to measuring complexity in quantum field theories that was motivated by holography. This approach arose in the context of tensor networks, which provide variational ansatze for many-body quantum wavefunctions. In particular, optimized tensor networks such as, e.g., MERA [17, 414] or its continuous counterpart cMERA [18, 98], give rise to a geometric structure that reflects the entanglement patterns of a quantum state. An analogy between the tensor network approach to constructing quantum states and the structure of the holographic duality had already been noted by Swingle [239]. This proposal seeded a new line of research that led Susskind [6] to argue that the growth of the volume of Einstein-Rosen bridges should be related to complexity, estimated by the number of tensors (the effective volume of the network) in the tensor-network representation of the time-evolved thermofield-double state [245]. In Sec. 8.1, we review a notion of complexity in CFT called path-integral complexity inspired by the parallelism with tensor networks. In Sec. 8.2 we discuss the holographic interpretation of this quantity in AdS spacetime.
@@ 8.1 Complexity from path-integral optimization
Motivated by these developments, Ref. [38] aimed at extending intuitions from the study of tensor networks to general, strongly-interacting QFTs and to use these ideas to define a natural measure of complexity using path integrals. A brief review of their construction follows.
In QFTs, wavefunctions can be prepared by performing path integrals over physical space and Euclidean time. They are computed by integrating over all field configurations, subject to specified boundary conditions. As an example, consider two-dimensional CFTs on a flat, Euclidean plane $\mathbb{R}^{2}$ with coordinates $(\tau,x)$, and collectively denote all fields in the CFT by $\Phi(\tau,x)$. The ground state wave function $\Psi_{\rm CFT}[\Phi(x)]$, which in QFTs is a functional of the configuration at the time slice $\tau=0$, $\Phi(x)\coloneqq\Phi(x,\tau=0)$, is defined as
@@
where $S_{\rm CFT}$ is the Euclidean action of the 2D CFT and $D\tilde{\Phi}$ is the path integral’s measure. [^39] Formally one may need to deform the theory by putting it in a finite volume or introducing a small mass so that the spectrum is gapped and the path integral (143) acts as a proper vacuum projector.
The main new idea in [38] was to introduce a non-trivial background metric on the space where the path integral (143) is performed (keeping the boundary conditions for the fields fixed), and to interpret this metric as an (unoptimized) density of gates of a continuous tensor network. For our 2D CFT example, all the metrics can be put into the Weyl-flat form parametrized by a scalar function $\phi(\tau,x)$ as follows:
@@
Figure 13: A cartoon explaining path-integral optimization in CFTs as an analog of tensor-network optimization. Figure from [38] with Euclidean time denoted as $z=-\tau$.
Next, we introduce the actual optimization procedure. Suppose we compare the wave functions computed with a background flat metric (denoted with $\Psi_{\text{CFT}}^{\phi_{0}}[\Phi(x)]$) and with a curved metric (denoted with $\Psi_{\text{CFT}}^{\phi}[\Phi(x)]$), they are proportional to each other
@@
where $c$ is the central charge of the 2D CFT. [^40] The coefficient $\mu$ of the potential term can be set to $1$ by shifting $\phi$. Intuitively, the proportionality factor in (145) can be interpreted as a measure of how many repetitions of path-integration per site we perform to construct the state. Hence, the optimal path integral is selected by minimizing the Liouville action with respect to $\phi$, i.e., picking the metric that solves the Liouville equation
@@
with the boundary condition (144). Recall that we are interpreting the metric as encoding the density of gates in the tensor network. This procedure is therefore called the path-integral optimization.
For example, for the CFT vacuum that is obtained by performing a Euclidean path integral on the half-plane, the optimal geometry is given by the hyperbolic plane. Similar results can be obtained for path integrals on the circle, that prepare the vacuum state of a CFT on the circle; as well as path integrals on a strip of size $\beta$ used to prepare the thermofield-double state with inverse temperature $\beta$. In the former case, the background obtained via the path-integral optimization procedure is the hyperbolic geometry of the Poincaré disc. In the latter case, the background geometry is the hyperbolic strip, also called the Euclidean trumpet geometry.
Similar optimization was also defined for the vacuum and thermal states of higher-dimensional CFTs [39], but with a prescription to optimize only over the Weyl factor of the background metric on which the CFT path integral is evaluated (unlike in 2D, this is only a subset of possible geometries in higher-dimensions). Again, it was found that optimal metrics have a constant negative Ricci scalar curvature i.e., are hyperbolic.
It is natural to think about these metrics as continuous counterparts of effective tensor-networks geometry that emerges after the optimization procedure in many-body systems. In fact, in the spirit of tensor network discussions in holographic CFTs, we can interpret the metrics from the optimization of path-integrals in $CFT_{d}$ as particular slices of their higher-dimensional $AdS_{d+1}$ dual. We will return to this point in Sec. 8.2.
Another idea from [38, 39], dubbed path-integral complexity, was to interpret the Liouville action as a novel measure of complexity for wave functions prepared using Euclidean path integrals. The main reason to advocate this interpretation was that the on-shell Liouville action is proportional to the volume of the space on which the path integral is computed. Indeed, after fixing the normalization such that the integral of the Weyl factor $e^{2\phi}$ over a unit area (a square region of size $\epsilon$ in the original metric) has a value 1 (see detailed discussion in [415]), we can interpret the volume of the Euclidean space as the number of tensors in the optimal network. The on-shell action associated with the CFT vacuum state reads
@@
This result is also consistent with qualitative expectations and holographic conjecture [25], which estimates the leading divergence in complexity of the CFT state by the spatial volume.
@@
Later on, researchers extended the path-integral optimization method to other settings. As we already mentioned, [38, 39] proposed a higher-dimensional generalization of the action (148). The the optimization of 2D CFTs deformed by relevant operators was studied in [415], where a modified Liouville action was minimized over both metrics and local deformation couplings. The authors of [416] investigated path integrals for inhomogeneous CFTs, that are usually interpreted as 2D CFTs on non-trivial background metrics. From the perspective of the optimization, the starting point is the CFT on a non-trivial background metric ($\hat{g}_{ab}$ instead of $\delta_{ab}$) and then consider $g_{ab}=e^{2\phi}\hat{g}_{ab}$ as the metric for optimization. This leads to additional background Ricci curvature term $\hat{R}$ in the Liouville equation which can nevertheless be solved using standard uniformization techniques.
Soon after [38], reference [417] argued that a variation of path integral complexity in holographic 2D CFTs leads to Einstein’s equations in the bulk of AdS3. Moreover, the Liouville action (or its covariant version, called the Polyakov action) motivated the treatment of Nielsen’s complexity by introducing trajectories on the unitary manifold generated by sequences of 2D conformal transformations [40] (see Sec. 5.6.2).
@@
The formulation in terms of a Liouville action also inspired a series of developments in the tensor-network community. They used the path integral approach to cMERA to define tensor networks in terms of the background geometry on which the state is prepared [418, 419]. In particular, they showed how the background metric parameters translate into details of the tensor network. This idea was later used in [420] to argue for a Nielsen-type interpretation of the Liouville action in terms of an appropriate choice of the cost function on the complexity geometry. As opposed to the developments on Nielsen’s complexity discussed in Sec. 5, the work [420] necessarily utilized circuits that contained both unitary and Hermitian gates associated with components of the energy-momentum tensor operator. This is similar in spirit to the results discussed in Sec. 7.5.1 of [370] and has the same origin: Euclidean time evolution is an efficient and natural method for preparing useful states. This points towards a mathematical generalization of the Nielsen’s complexity that allows for exponentiation of anti-Hermitian (giving rise to unitary transformations), as well as Hermitian matrices (giving rise in particular to Euclidean time evolution).
@@
Finally, the results in Ref. [420] pointed out that a general cost function for path integral optimization will also contain positive powers of the cut-off $\epsilon$ associated with higher derivatives of $\phi$ in the spirit of an effective field theory. However, the solutions of Eq. (147) acquire derivatives that are themselves of the order of $1/\epsilon$, breaking the hierarchy within these more general cost functions that left the contribution (146) as the dominant term. Thus we need a more systematic inclusion of the UV cut-off $\epsilon$, bringing us to the subject of the next section.